OpenStudy (anonymous):

The diagonal AC is drawn in parrollelogram ABCD. Which Method can not be used to prove that abc~cda? a. SAS b.SSS c. ASA d.SSA PLEASE SOMEONE HELP!

OpenStudy (anonymous):

SAS

OpenStudy (anonymous):

@Noemi95 thank you so much. would you mind helping me sommore?

OpenStudy (anonymous):

No problem dime, and sure. I don't ensure knowing everything, but I can try!(:

OpenStudy (anonymous):

@Noemi95 you know what they say the more the merrier. :) lol In circle O, a diameteer has endpoints (-5,4) AND (3,-6). What is the length of the diameter?

OpenStudy (anonymous):

2* sqrt 41

OpenStudy (anonymous):

Draw a right angle triangle ABC where A is point (-5,4) , C is point (3-6) and B is the point that makes a right angle with the other two points (3,4) The diameter of the circle is AC and is the hypotenuse of triangle ABC AB = distance along the x axis ie -5 to + 3, ie distance AB = 8 BC = distance along the y axis ie 4 to -6 ie distance = 10 Therefore using Pythagoras's theorem AC^2 = AB^2 + BC^2 AC^2 = 64 + 100 AC^2 = 164 AC = square root 164 AC = square root (4*41) as square root 4 = 2 AC = 2 square root 41

OpenStudy (anonymous):

@Noemi95 If AB=5x-9, DC=x+3 and EF = 2x+2 what is the length of EF??

OpenStudy (anonymous):

Well, let's see. To figure out the value of x, we need to add of the sides then solve for x. So, 5x-9+x+3+2x+2= 8x-4 (now solve for x) P=8x-4 +4 +4 4/8=8x/8 x=0.5 Now plug in this value for EF. 2(0.5)+2=3 "EF=3"